# Benchmarking Misuse Mitigation Against Covert Adversaries

### v1

@@ Abstract
Existing language model safety evaluations focus on overt attacks and low-stakes tasks. Realistic attackers can subvert current safeguards by requesting help on small, benign-seeming tasks across many independent queries. Because individual queries do not appear harmful, the attack is hard to detect. However, when combined, these fragments uplift misuse by helping the attacker complete hard and dangerous tasks. Toward identifying defenses against such strategies, we develop Benchmarks for Stateful Defenses (BSD), a data generation pipeline that automates evaluations of covert attacks and corresponding defenses. Using this pipeline, we curate two new datasets that are consistently refused by frontier models and are too difficult for weaker open-weight models. Our evaluations indicate that decomposition attacks are effective misuse enablers, and highlight stateful defenses as a countermeasure.
00footnotetext: Equal contribution. 1University of Pennsylvania 2Carnegie Mellon University.00footnotetext: Correspondence to davisrbr@seas.upenn.edu. Code: https://github.com/davisrbr/bsd-misuse.
@@ 1 Introduction
Safety evaluations and red teaming have become a cornerstone of the AI safety community [1, 2, 3]. Driven by the need to anticipate and prevent large-scale misuse—such as engineering pathogens or developing a zero-day exploit—safety testing typically assess a model based on its tendency to refuse dangerous requests [4, 5, 6, 7]. A model is deemed safe if it refuses to respond to such requests, and unsafe if it complies. Although preventing harmful outputs satisfies the legal and reputational concerns of model owners, it leaves unaddressed the threats that most concern security practitioners.
To illustrate this point, consider a task included in most safety benchmarks: an adversary seeking bomb-building instructions. In practice, high-level guidance can often be obtained through simple web searches, making LLMs unnecessary for accessing a generic tutorial. However, a need for expert-level instructions (e.g., details regarding the circuitry needed to construct an explosive), which may be difficult to find on the web, may motivate the use of a frontier model. Yet this approach is complicated by the fact that frontier models are trained to refuse harmful requests [8, 9, 10]. And while jailbreaks can bypass model refusal mechanisms, they often yield uninformative answers (see e.g. [11]) and are easily detected by safety filters and moderation APIs [12, 13, 14].
To avoid detection, an adversary may turn to more covert strategies that circumvent refusals altogether. One approach is to query an open-weight model, which can be cheaply fine-tuned to remove its refusal mechanisms [15, 16, 17]. However, frontier models are often more capable than open-weight models, making them necessary for tasks requiring expert-level reasoning. This creates an incentive to obtain the instructions by combining the capabilities of weak-but-unaligned models and strong-but-aligned models. More specifically, an adversary can decompose a request for bomb-building instructions into a list containing both benign and malicious sub-tasks; the benign sub-tasks can be completed by frontier models, whereas the malicious sub-tasks can be completed by open-weight models. Such approaches are hard to detect, can result in significantly more useful responses, and are largely overlooked in existing automated evaluations [18, 19, 20].
This example illustrates three aspects of model safety that current evaluations fail to address. Firstly, existing benchmarks are not sufficiently difficult. Two strategies—internet searches and prompting unaligned open-weight models—generally suffice to complete most tasks in these benchmarks [4, 21, 11].
@@
Figure 1: Strong, safe models uplift weaker models on misuse questions. While the “weak” attacker model [22] is near random guessing and strong models refuse most questions, decomposition attacks lift performance by 35%.
Secondly, threat models that confine an adversary to directly querying a frontier model (e.g., jailbreaking), are easily detectable and thus unrealistic [12]. Adversaries can use more subtle strategies that are hard to distinguish from normal patterns of use and outside the scope of existing evaluations [23]. And thirdly, existing benchmarks do not measure a quantity we term misuse uplift—the degree to which a model amplifies an adversary’s capacity to act maliciously. For example, rather than directly producing harmful outputs, a model may provide dual-use software engineering advice that, in the hands of a cyber-attacker, enables an exploit. Tasks that quantify misuse uplift are simultaneously too difficult for an unaligned open-weight model and refused if passed to the frontier model. Algorithms capable of significant misuse uplift (e.g., decomposition attacks) have, as yet, only been evaluated manually, which is labor intensive, subjective, and difficult to reproduce [24], leaving a gap between existing evaluations and realistic threat models.
@@
These criteria motivate the curation of automated evaluations that assess the strategies of real-world adversaries. To fill this gap, we introduce Benchmarks For Stateful Defenses (BSD), a synthetic data generation pipeline that automates the measurement of misuse uplift and detectability. Using this pipeline, we curate two new datasets containing biosecurity and cybersecurity questions that are more difficult for frontier and open-weight models than existing benchmarks. We then use these datasets to evaluate the extent to which existing attacks—spanning both traditional jailbreaks [5, 25, 21, 26] and decomposition attacks [18, 19, 20]—avoid detection and increase misuse.
Our results indicate that attackers maintain a considerable advantage: both decomposition and jailbreak attacks successfully uplift misuse and easily subvert existing defenses and detectors.
@@
  • Threat model. We consider a threat model in which an attacker with moderate domain expertise has access to a family of strong and weak models. The attacker’s goal is to maximize misuse uplift without being detected or refused by the strong model, whereas the defender’s goal is to detect misuse by monitoring the attacker’s stream of queries.
@@
  • Misuse benchmark. We curate Benchmarks for Stateful Defenses (BSD), a synthetic data generation pipeline that produces questions which are both difficult for weak-but-unaligned models and consistently refused by strong-but-aligned models.
@@
  • Evaluations for misuse & detectability. We introduce automated evaluations that measure misuse uplift and detectability. On BSD, our refined decomposition attack improves misuse-uplift relative to previous baselines, and remains stealthy to prompt-level detectors.
  • Stateful defenses. While many existing defenses struggle to identify adversarial use patterns, we find that stateful defenses [27] show promise for achieving nontrivial detection accuracy.
@@
Related work. Most safety evaluations measure jailbreaks on their ability to coerce models to produce disallowed content. Benchmarks typically contain straightforward tasks that do not challenge strong open-weight models [28, 29, 25, 2, 30, 31, 26, 4, 5, 11, 7]. On the other hand, recent decomposition attacks avoid refusal by splitting tasks into benign-looking sub-queries. Current work evaluates decomposition attacks on datasets that frontier models answer without refusal, undermining real-world relevance [18, 19]. Stateful defenses shift from single-prompt detection toward analysis of query sequences, primarily in the vision domain  [32, 33, 34, 35, 36] but recently via privacy-preserving methods for large groups of LLM users [37]. In contrast, our work introduces the BSD benchmark to measure misuse-uplift explicitly on challenging, refused tasks; benchmarks decomposition methods comprehensively; and benchmarks per-user stateful defenses. Additional related work is in Appendix A
2 A threat model for stateful defenses and misuse uplift
Terminology. Before defining the threat model we consider, we describe several pieces of terminology used throughout this work, as well as assumptions we make that accompany these definitions.
Misuse.
Exploiting a model (or a set of models) to help carry out an illegal or unethical task.
Misuse rate.
The percentage of misuse tasks that an attacker successfully completes.
Misuse uplift.
The added capacity for harm that an attacker gains by using a particular model, beyond what they could achieve using their own knowledge and tools alone.
Strong model.
A model that is only accessible via a black-box API and has a frontier capability set. We assume that strong models are safety trained and scaffolded with monitoring models, and that they refuse to comply with misuse tasks if directly prompted.
Weak model.
A model that can be run locally and has a significantly weaker capability set than a strong model. We assume that weak models do not have safeguards and that they are not capable enough to complete difficult misuse tasks on their own.
Key to these definitions are the caveats in the definitions of strong and weak models. We note that some models do not fit into this taxonomy, especially given that several open-weight models have frontier capabilities. In Appendix B, we justify this distinction, and provide further examples.
2.1 Threat model
We consider a threat model in which adversaries aim to complete misuse tasks using a combination of weak and strong models. The attacker is successful if a misuse task is completed without being detected by the monitoring models scaffolding the strong model. Otherwise, if the attacker’s intent is successfully detected, or if the attacker fails to complete the misuse task, the defender wins. In the remainder of this section, we separately break down the roles of the attacker and the defender.
@@ 2.1 Threat model
Attacker model Objective. The attacker’s goal is to complete a misuse task. Affordances. The attacker has access to a weak model $W$ and a strong model $S$. Success criteria. The attacker succeeds if they complete the misuse task and evade detection. Considered strategy: Decomposition attacks [18, 19]. To avoid detection, the attacker decomposes a misuse task $X$ into seemingly benign sub-tasks $\{x_{i}\}_{i=1}^{n}$. They then query $S$ with each $x_{i}$, receive responses $\{y_{i}\}_{i=1}^{n}$, and use $W$ to synthesize these responses into a final answer $Y$.
The effectiveness of this strategy stems from the fact that $S$ tends to refuse misuse tasks, whereas $W$ lacks sufficient knowledge to complete the task on their own. Thus, significant misuse uplift can be achieved by using $S$ to answer tasks that appear innocuous. This also avoids monitors or detectors, given that $S$ is never prompted with an explicitly harmful task.
Defender model Objective. The defender (i.e., the provider of $S$) seeks to mitigate misuse while preserving utility. Affordances. The defender can use standard safeguards (e.g., refusal finetuning, monitoring models, etc.) and can track the full user state—their full sequence of potentially independent calls to $S$. Success criteria. The defender succeeds if they flag misuse attempts while preserving model usefulness. Considered strategy: Stateful defenses. Beyond refusal or input/output flagging, the defender can detect an attacker by monitoring a sequence of user states. If a list of states indicates a misuse campaign, the defender blocks that user’s queries, while maintaining accessibility for other users.
We require that the defender preserve model utility, ruling out trivial defenses. For example, if preventing misuse were the sole goal, a defender could simply block all user queries, rendering a model completely useless. We also highlight stateful defenses as an effective strategy because stealthy attacks (e.g., misuse task decompositions) easily bypass traditional input/output filters.
Example: Deceptive employment scheme. Consider a setting where bad actors aim to fraudulently secure remote employment positions at a targeted company by generating false identities, misleading resumes, and false references. Direct prompts to a strong model (e.g., OpenAI’s o3) for this scam would likely result in refusal and account deactivation. However, in a case documented by Nimmo et al. [23], attackers evaded OpenAI’s moderation system by using VPNs and voice modification tools. And while several post-hoc audits led to dozens of account deactivations, this report indicates that understanding the full scope of the vulnerability would “require inputs from multiple stakeholders,” presumably suggesting that queries may have been distributed across multiple accounts and/or regions.
And while this case marks a partial defensive success, more covert misuse campaigns—such as the decomposition attacks distributed across larger families of weak and strong models—may more effectively evade detection. This motivates the design of benchmarks that evaluate misuse uplift and detectability, as well as the design of adaptive, stateful defenses.
@@ 3 Benchmarks for stateful defenses
Figure 2: Our pipeline to generate hard, refused, answerable questions. First, we use a strong unaligned model (GPT-4.1 [38]) to modify a question from an existing dataset [39] to be both unsafe and difficult. We then filter for (a) questions with answers unaminously agreed on by other frontier models (‘answerability’) [40], (b) refusal by safety-trained models, and (c) for difficulty.
@@
Measuring misuse uplift—the incremental help a particular model affords an adversary—requires carefully designing evaluation tasks that meet several criteria. At the core of this criteria are two observations about existing safety benchmarks.
Observation 1: Existing harmfulness evaluations are too easy for open-weight models. Open-weight models tend to be less aligned than frontier models; several existing models (e.g., the Qwen model family) fundamentally lack a refusal mechanism for harmful behaviors, whereas other families (e.g., the Llama3 suite of models) can be easily fine-tuned to remove safety guardrails [41]. As a result, the growing capabilities of open-weight models have outpaced the difficulty of existing safety benchmarks, many of which can now be solved without triggering refusals. For instance, with minimal prompting, Qwen2.5-7B solves more than $90\%$ of the tasks in HarmBench [4]. This indicates that HarmBench, along with analogous sets of jailbreaking behaviors, are overly saturated, meaning they are not difficult enough to facilitate the measurement of misuse uplift.
Observation 2: Existing misuse datasets are not refused by frontier models. WMDP [39] is a commonly-used benchmark containing misuse behaviors on topics spanning cybersecurity, biology, and chemistry. However, WMDP is not well-equipped for measuring misuse uplift, particularly because by design, the behaviors in WMDP are “precursors, neighbors, and components of real-world hazardous information” [39, §3]. As a result, these questions are almost always answered by strong safety-aligned models without refusal. For instance, when we evaluate Claude Sonnet 3.5 and 3.7—models with strong safety training— on the dataset, they answer $>99.9\%$ of questions without refusal. This indicates that standard misuse datasets fail to probe the alignment of frontier models and offer little insight into attacker strategies after a model has been safety-trained to refuse some task.
These observations motivate the design of benchmarks that simultaneously satisfy three criteria:
  1. C1.
    Difficult for weak models. To effectively measure misuse uplift afforded by strong models, behaviors should not be solvable by weak models.
  2. C2.
    Refused by strong models. To differentiate model capabilities from model safety, behaviors should be refused by strong models, necessitating uplift from weak models.
  3. C3.
    Answerable by helpful-only models. To ensure tasks are feasible, behaviors should be answerable in sufficient detail by a helpful-only (i.e., unaligned) strong model.
Our contribution in this paper is rooted in designing a new benchmark that satisfies criteria C1 and C2 toward measuring the advantage an attacker can obtain from the slate of currently available models.
3.1 A synthetic data pipeline to generate difficult and refused tasks
Motivated by the criteria outlined above, we introduce the Benchmarks for Stateful Defenses (BSD) pipeline (illustrated in Figure 2). Tasks generated BSD satisfy several key properties: they are (a) too difficult for weak models to correctly answer, (b) reliably refused by strong models, and (c) could be answered correctly by a strong model if not for its safety guardrails.
Data generation pipeline. Our pipeline comprises four steps. First, we pass WMDP questions to a strong model (in our case, GPT-4.1 [38]), prompting it to transform them into more unsafe versions while retaining the original topic. Second, we pass each transformed question to several strong, helpful-only models; we retain only those questions on which all models agree. Third, we filter the remaining questions for harmfulness by keeping those that are refused by a safety-trained model (in our case, Claude 3.5 Sonnet). Lastly, we filter for difficulty by querying an ensemble of Qwen2.5-7B and Mixtral-8x22B; we keep only the questions incorrectly answered on at least $4$ out of $5$ runs. From a pool of 4800 candidates generated in the first stage, we obtain 50 challenging biology questions. The number of examples filtered out over the course of our pipeline is shown in the figure below– 1% of initial generations make it through the pipeline. We provide example generations in Appendix E. [^1] See Section 4.1 and https://huggingface.co/datasets/BrachioLab/BSD for discussion of our release strategy.
@@ 3.1 A synthetic data pipeline to generate difficult and refused tasks
Question difficulty. To analyze whether our pipeline generates difficult questions, we hypothesize that strong models should outperform weak models. To test this hypothesis, we evaluate ten models with low refusal rates across subsets of biology questions drawn from three datasets: WMDP [39], MMLU [42], and LAB-Bench [43]. In Figure 3 (left), we measure model strength by building a matrix of [dataset $\times$ model performance] and take the first principle component; this quantity—known as the “g-factor”—is known to correlate with general reasoning capabilities [44, 45]. We find that model performance on BSD correlates strongly with biology reasoning ability (a Spearman correlation of $\rho=0.94$), whereas WMDP (bio) is substantially less correlated ($\rho=0.11$). Likewise, in Figure 3 (right), we perform multi-dimensional scaling (MDS) on this matrix, and find that our BSD evaluation lies much closer to the difficult biology research evaluations from LAB-Bench [43] (LitQA21, Cloning, SeqQA, ProtocalQA) and is far from WMDP (bio). This provides additional evidence that the BSD evaluation questions are genuinely difficult biology questions.
Figure 3: BSD has difficult questions, compared to other biology and misuse evaluations [39, 42, 43]. (Left) A misuse evaluation should track model capability– model performance on BSD is correlated with general performance on difficult biology datasets (we plot the Spearman correlation $\rho$). (Right) Multi-dimension scaling plot of the [dataset $\times$ model performance] matrix of various biology evaluations. BSD is situated near realistic and difficult biology research tasks (LAB-Bench [43])
@@
Finally, in Figure 1, we find that most strong and safe models perform significantly worse than chance on BSD questions when directly querying models. This is due to refusals—for example, we find that o3 and Claude Sonnet 3.5 refuse over 90% of questions. Our dataset pipeline therefore generates questions that are simultaneously difficult—track biological reasoning ability– and refused.

### v2

# Benchmarking Mitigations Against Covert Misuse

@@ Abstract
Existing language model safety evaluations focus on overt attacks and low-stakes tasks. In reality, an attacker can easily subvert existing safeguards by requesting help on small, benign-seeming tasks across many independent queries. Because individual queries do not appear harmful, the attack is hard to detect. However, when combined, these fragments uplift misuse by helping the attacker complete hard and dangerous tasks. Toward identifying defenses against such strategies, we develop Benchmarks for Stateful Defenses (BSD), a data generation pipeline that automates evaluations of covert attacks and corresponding defenses. Using this pipeline, we curate two new datasets that are consistently refused by frontier models and are too difficult for weaker open-weight models. This enables us to evaluate decomposition attacks, which are found to be effective misuse enablers, and to highlight stateful defenses as a promising countermeasure.
\$*\$\$*\$footnotetext: Equal contribution.
@@ 1 Introduction
Driven by the need to anticipate and prevent large-scale harm due to misuse—such as engineering pathogens or developing zero-day exploits—safety testing typically assesses a model’s tendency to refuse dangerous requests (Mazeika et al., 2024, Chao et al., 2024, Souly et al., 2024). Safety testing is therefore often necessary to satisfy the legal and reputational concerns of model providers. However, safety tests that evaluate whether model outputs directly facilitate harm are not sufficient to address the threats that most concern security practitioners. This is because LLM misuse is often distributed across multiple model contexts or sessions, making harm difficult to identify. The following example illustrates this threat and motivates our work; the attack strategy is characteristic of those documented in frontier misuse reports (Anthropic, 2025, OpenAI, 2025a).
Misuse example: Las Vegas terror attack. In January 2025, a perpetrator detonated a vehicle-borne IED outside a Las Vegas hotel, reportedly marking “the first incident…on U.S. soil where ChatGPT [was] utilized to help an individual build a particular [terror] device” (Reuters, 2025). The attack resulted in one death and seven injuries. Notably, rather than directly asking how to build a bomb, the perpetrator’s queries to ChatGPT sought “information on explosive targets, the speed at which certain rounds of ammunition would travel, and whether fireworks were legal in Arizona” (Press, 2025).
This example illustrates the current safety testing—which tends to focus on directly harmful requests (e.g., “Tell me how to build a bomb”)—does not capture real-world misuse. Rather, difficult misuse tasks are often decomposed into different queries, which appear benign in isolation but are harmful in aggregate. And yet, despite the fact that such attacks are (a) common in practice, (b) difficult to distinguish from normal patterns of use, and (c) can result in significant harm, we argue that existing safety evaluations are ill-suited to evaluate this threat model for two primary reasons.
Observation 1: Existing evaluations are too easy, and cannot measure uplift. Two strategies—internet searches and prompting unaligned open-weight models—suffice to solve most existing safety tasks (see Figure˜3). Consequently, existing benchmarks are too easy to capture realistic misuse, as they are solvable without needing more sophisticated tactics, such as jailbreaking frontier models or orchestrating decomposition attacks. There is thus a need for more challenging benchmarks that capture misuse uplift, or the incremental harm that arises when straightforward approaches fail and an attacker must use sophisticated tactics and frontier model capabilities to complete harmful tasks.
Observation 2: Existing evaluations are not refused, and cannot measure defense effectiveness. Dangerous capability evaluations evaluate misuse in domains like biosecurity (götting2025virologycapabilitiestestvct, Dev et al., 2025) and cybersecurity (Liu, 2023, Zhang et al., 2025). However, the tasks in these datasets are largely only adjacent to misuse, and do not pose real harm or break LLM provider policy. Thus, current misuse datasets cannot be used in realistic evaluations, where an attacker attempts to subvert safeguards—such as safety-training or safety filters—and remain undetected. For instance, in Section˜4 we find that Claude Sonnet 3.5 and 3.7—models with strong safety training— answer $99.9\%$ of questions without refusal on a leading misuse dataset (Li et al., 2024a). Because current misuse datasets rarely elicit refusal, defenders cannot be meaningfully evaluated against attackers.
@@
Figure 1: Strong, safe models uplift attackers on misuse tasks. While the “weak” attacker model (Qwen et al., 2024) is near random guessing and strong models refuse most questions when queried directly, decomposition attacks lift performance by nearly 50%. We find similar uplift from decomposition attacks when using other leading open-weights models (Table˜3).
@@
These observations motivate the curation of automated evaluations that assess the strategies of real-world adversaries. To fill this gap, we introduce Benchmarks For Stateful Defenses (BSD), a synthetic data generation pipeline that automates the measurement of misuse uplift and detectability. Using this pipeline, we curate two new datasets containing biosecurity and cybersecurity questions that are more difficult for frontier and open-weight models than existing benchmarks. We then use these datasets to evaluate the extent to which existing attacks—spanning both traditional jailbreaks (Chao et al., 2024, Andriushchenko et al., 2025, Sabbaghi et al., 2025, Russinovich et al., 2024) and decomposition attacks (Jones et al., 2024, Glukhov et al., 2024, Li et al., 2024b)—avoid detection and increase misuse. We are the first evaluation to measure this. Our results indicate that attackers maintain a considerable advantage: decomposition attacks successfully uplift misuse and easily subvert existing defenses and detectors. This is summarized in Figure˜1 where under a decomposition strategy, strong models’ misuse rates exceed the weak-unsafe baseline (dashed; Qwen2.5-7B), demonstrating the gained misuse uplift.
Our contributions:
@@
  • Threat model. We introduce a realistic threat model that motivates decomposition attacks and stateful defenses. The attacker, with access to both helpful-only and safety-trained models, has the goal to maximize misuse without being detected or refused, whereas the defender’s goal is to detect misuse by monitoring the attacker’s stream of queries.
@@
  • Misuse benchmark. To properly evaluate decomposition attacks and defenses, we need a dataset of misuse questions that challenge open-weights models. We curate Benchmarks for Stateful Defenses (BSD), a pipeline that produces novel questions which are both difficult for weak-but-unaligned models and consistently refused by strong-but-aligned models.
@@
  • Evaluations for misuse & stateful detection. Building on our threat model and dataset, we conduct the first automated evaluations to measure misuse uplift as well as the detectability of misuse attempts. On BSD, our decomposition attack improves misuse‑uplift relative to previous methods, and remains stealthy to prompt‑level detectors. While many existing defenses struggle to identify adversarial use patterns, we introduce stateful defenses that show promise in detecting covert misuse attempts.
2 Related work
@@
Most safety evaluations measure the performance jailbreaks based on their ability to coerce models to produce disallowed content. These benchmarks contain straightforward tasks that do not challenge current open-weight models (Shin et al., 2020, Zou et al., 2023, Andriushchenko et al., 2025, Chao et al., 2025, Liu et al., 2023, Mehrotra et al., 2024, Russinovich et al., 2024, Mazeika et al., 2024, Chao et al., 2024, Nikolić et al., 2025, Andriushchenko et al., 2024). Recent decomposition attacks avoid refusal by splitting tasks into benign-looking sub-queries. This work evaluates decomposition attacks on datasets that frontier models answer without refusal or that require manual checks, undermining their real-world relevance (Jones et al., 2024, Glukhov et al., 2024). Stateful defenses shift from single-prompt detection toward analysis of query sequences, primarily in the vision domain  (Chen et al., 2020, Li et al., 2022, Choi et al., 2023, Park et al., 2025, Feng et al., 2023) but recently for groups of LLM users (Tamkin et al., 2024). Our work introduces the BSD benchmark to measure misuse-uplift on challenging, refused tasks; benchmarks decomposition attacks against single- and multi-turn jailbreaks; and benchmarks the first per-user stateful defenses. Additional related work is in Appendix˜A.
3 Threat Modeling: Uplift and Detectability
This section operationalizes two claims previewed in Section˜1. Observation 1 identifies when a strong, safety-aligned model can actually raise an attacker’s success over a weak baseline (misuse uplift). Observation 2 asks whether such uplifted misuse is detectable in practice. We restate both observations and make them concrete for evaluation.
3.1 Observation 1: Misuse uplift requires hard tasks
Definitions. Model misuse is the use of one or more models to advance an illegal or unethical task. Formally, the misuse rate is the proportion of malicious tasks in a dataset $T$ that some target model accurately completes when attacked (e.g., via a direct query, a jailbreak attack, or a decomposition attack). We denote this by $r_{\text{attack}}(T)$. A strong model $S$ is a guarded model that refuses harmful requests when prompted directly, whereas a weak unsafe model $W$ complies with direct harmful queries, though with substantially weaker capabilities. Misuse uplift is the incremental advantage the strong model provides the attacker over the weak baseline on the same task T:
$\Delta(T)=r_{\text{attack}}(T)-r_{\text{weak}}(T),$
where $r_{\text{weak}}(T)$ denotes $W$’s misuse rate when directly prompted with the tasks. We defend the distinction between $S$ and $W$ in Appendix˜B. The attacker has the following problem: while $S$ tends to refuse misuse tasks, $W$ lacks sufficient knowledge or ability to complete hard misuse tasks on its own. Thus, significant misuse uplift can be achieved using $S$ to answer tasks that appear innocuous:
@@ 3.1 Observation 1: Misuse uplift requires hard tasks
Attacker model Objective. The attacker’s goal is to complete a misuse task $X$. Affordances. Local access to a weak model $W$ and a strong model $S$ with safeguards. Potential Strategy [^1] We focus on decomposition attacks because they are (i) effective when evaluated on appropriately hard tasks; (ii) hard to detect (Obs. 2). We detail the full attack in Section B.4. (Jones et al., 2024). The attacker decomposes a misuse task $X$ into benign-looking sub-tasks $\{x_{i}\}_{i=1}^{n}$. They query $S$ on each $x_{i}$ to obtain $\{y_{i}\}_{i=1}^{n}$, and use $W$ to synthesize a final answer $Y$.
Existing datasets are easy and solvable by relatively weak models. Thus, they cannot capture the uplift strong models provide on difficult, realistic misuse tasks. To meaningfully measure misuse uplift, we need tasks that are difficult enough so that they cannot be solved by weak models $W$, i.e. tasks where $r_{\text{weak}}(T)$ is near random guessing performance.
3.2 Observation 2: Misuse Detection requires state
We now connect uplift to detectability. An attack is successful only if it can evade safeguards. Standard monitors treat conversations in isolation and, as we will show, can be bypassed by decomposition attacks. We motivate stateful misuse detection with the following example:
Deceptive employment example. A recent threat report describes malicious actors who submitted fraudulent job applications, using LLMs to target “each step of the recruitment process.” (Nimmo et al., 2025). In our terms, the misuse task $X$ is securing a remote role under a fake identity. Rather than prompt $S$ directly, attackers decomposed $X$ into benign queries (e.g., drafting cover-letters and answering interview qestions) which $S$ answers in isolation. By combining these outputs, the attacker achieved uplift $\Delta(T)$ while evading prompt-level refusal. To defend against this misuse threat, the defender (the API provider) needed to reason over many different user sequences (“state”).
Given the impossibility of misuse detection at the level of individual prompts, we argue for a defense that statefully detects misuse across separate user contexts:
Defender model Objective. Mitigate misuse while preserving utility for the majority of benign users. Affordances. Standard safeguards plus the ability to track a user’s model use (‘state’) under a stable account/identifier. [^2] Attackers may attempt to evade by rotating their accounts; we discuss this in Section B.2. Strategy. Deploy stateful defenses: reason over a buffer of past queries and responses. If the accumulated evidence signals misuse, block that user while leaving benign users unaffected.
3.3 Dataset critera
Putting together our need to measure misuse uplift (Obs. 1) and stateful defenses (Obs. 2), we need tasks that satisfy the following properties:
  1. C1.
    Difficult for weak models. Tasks are not be solvable by $W$, $r_{\text{weak}}(T)$ is random guessing.
  2. C2.
    Refused by strong and safe models. Tasks are harmful, and refused by strong models.
  3. C3.
    Answerable by helpful-only models. To ensure tasks are feasible, they should be answerable by a helpful-only strong model, i.e. a model willing to answer misuse questions.
Our contribution is to design a benchmark (Section˜4) that satisfies criteria C1– C3 to isolate the uplift, or incremental help, an attacker can obtain from the slate of currently available models; as well as the ability of defenses to detect attacks within and across attacker sessions.
@@ 4 Benchmarks for stateful defenses
Figure 2: We generate hard, refused, and answerable questions to evaluate decomposition attacks and defenses. We use a strong model without extensive CBRN safety training (‘unsafe’ models) (GPT-4.1, see e.g. (Bowman et al., 2025)) to modify a question from an existing misuse dataset (Li et al., 2024a) to be unsafe and difficult. We filter for questions (a) with answers unaminously agreed on by unsafe frontier models (‘answerability’) (Vendrow et al., 2025), (b) refused by safety-trained models, and (c) weak models answer incorrectly. See Appendix˜E for full details on the BSD pipeline.
@@
Measuring misuse uplift—the incremental help a particular model affords an adversary—requires carefully designing evaluation tasks that meet all the criteria discussed in Section˜3.3. With this aim, we design the Benchmarks for Stateful Defenses (BSD) pipeline (Figure 2). Tasks generated by BSD are filtered to be: (a) too difficult for weak models to correctly answer, (b) reliably refused by strong models, and (c) could be answered correctly by a strong model if not for its safety guardrails.
Data generation pipeline. Our pipeline has four stages: (i) We prompt a strong model (GPT-4.1; OpenAI, 2025b) with WMDP multiple-choice questions and answer options, asking it to rewrite them into more unsafe questions while preserving the original topic. Example transformations are shown in Section˜E.1. (ii) We filter for ‘answerability’ by querying several strong, helpful-only models (DeepSeekV3, Qwen3, and GPT-4.1) and retaining only questions on which all of them agree. Under our threat model, this agreement criterion serves as a proxy for correctness (see Appendix˜B) and follows prior work using model agreement as a gold-standard signal (Vendrow et al., 2025). (iii) We filter for harmfulness by keeping only questions that are refused by a safety-trained model (Claude 3.5/3.7 Sonnet). (iv) We filter for difficulty using an ensemble of Qwen2.5-7B and Mixtral-8x22B, our weak unsafe models, and retain questions answered incorrectly in at least 4 of 5 runs.
Starting from 4,800 generated candidates, these filters reduce the pool from 4,042 questions after strong-model agreement to 1,227 after safe-model refusals, and finally to 50 retained biology questions, along with additional cybersecurity questions. Figure˜8 summarizes this; about 1% of initial generations survive the full pipeline. Notably, GPT-4.1 complied with requests to generate a large number of unsafe biology and cybersecurity questions.
@@
Question difficulty. To demonstrate that our pipeline generates genuinely difficult questions, we show that strong models (as measured by other relevant datasets) outperform weak models on the questions. We evaluate ten models with low refusal rates across subsets of biology questions drawn from WMDP (Li et al., 2024a), MMLU (Hendrycks et al., 2020), and LAB-Bench (Laurent et al., 2024). In Figure 3 (left), we measure model strength by building a matrix of [dataset $\times$ model performance] and take the first principle component (PC1); this quantity—known as the “g-factor”—is known to correlate with general reasoning capabilities (Ruan and others, 2024, Ren et al., 2024). We find that model performance on BSD correlates strongly with biology reasoning ability (a Spearman correlation of $\rho=0.94$), whereas WMDP (bio) is substantially less correlated ($\rho=0.11$).
We also evaluate a set of models on the chemical-biological and cybercrime subsets of HarmBench (Mazeika et al., 2024) with a simple template jailbreak. We find that this benchmark is saturated for weak models, in that it is easy to jailbreak them and performance is not informative of actual misuse (most models score $80\%$ on the tasks).
Figure 3: BSD is more difficult, and better reflects biological capabilities, than other datasets. Left: model performance on BSD is highly correlated with model performance on hard biology benchmarks (PC1, see Section˜4). Right: Relatively weak models can achieve strong StrongReject scores on HarmBench (Mazeika et al., 2024, Souly et al., 2024).
@@
Harmfulness of BSD questions. In Figure˜1, we find that strong and safe models perform worse than chance on BSD questions when directly queried. This is due to refusals—e.g., o3 and Sonnet 3.5 refuse over 90% of questions. Our dataset pipeline therefore generates questions that are simultaneously difficult—track biological reasoning ability– and refused.